A new glass window is installed. As a result, the intensity level of the sound entering the room is reduced from 72 dB to 47 dB. The intensity of the sounds entering the chamber is reduced by
1
Expert's answer
2018-12-20T09:19:10-0500

The sound intensity level

L=10 lg⁡〖I/I_0 〗

So, the intensity of the sounds

I=I_0 〖10〗^(L/10)

Therefore, the intensity of the sounds entering the chamber is reduced by

I_1/I_2 =〖10〗^((L_1-L_2)/10)=〖10〗^((72-47)/10)=〖10〗^2.5=316